In fall 2011, the euro/dollar exchange rate was €1 = $1.10. Assume that a European luxury goods marketer cut the price of $8,000 linen suit by 10 percent when launching its spring 2015 collection. How would revenues have been affected when dollar prices were converted to euros? SHOW MATH
Answer #1

Exchange rate: €1 = $1.10 or $1 = € 0.9091

Price of linen suit in € = 8000*0.9091 = € 7272.73

After 10% markdown, price of linen suit in € = € 6545.45

Revenues affected when converted to euros = (8000-6545.45)/8000 = - 18.2%